# Now at ]Max-Planck-Institut für Gravitationsphysik (Albert-Einstein-Institut) and Leibniz Universität Hannover, 30167 Hannover, Germany Now at ]Max-Planck-Institut für Gravitationsphysik (Albert-Einstein-Institut) and Leibniz Universität Hannover, 30167 Hannover, Germany

### v1

@@ IV.2.1 Open Shutter Power
To find $T_{\!{}_{\rm sci}}$, the ratio of power incident on the RC cavity mirror versus the power incident on the science detector was calculated. This was done several times before, during, and after the first science campaign and gave values between 80% and 90%. The value used for calibration was $85\%\pm 5\%$. The losses were mostly due to optical pickoffs used to sample the light reflected from the cavity for the PDH frequency stabilization system as well as other diagnostic purposes. Changes in the total reflectivity of the input optics that appear to be humidity dependent were also observed, along with clipping losses on the free aperture of the input optics after the beam expanding telescope in NR. Both of these effects played a role in the uncertainty of the results of $T_{\!{}_{\rm sci}}$.
IV.2.2 Central Optical Bench Transmissivity
Ex-situ measurements of the COB optics on the optical axis of the system were performed after the completion of the first science campaign (see Figure 6) with the COB removed from the vacuum system. As the transmissivities of these optics were strongly dependent on the angle of incidence ($\sim$ 1.5 ppm per degree for the mirrors sealing the light-tight housing and $\sim$ 150 ppm per degree for the MZ mirrors) the flat cavity mirror of the RC was used as an alignment reference to ensure that the laser used in the setup passed through the optics at the same angle as the HPL light during the open shutter periods of the science runs. The measurement procedure involved shining a laser coupled to an optical fiber on the mirror being tested. The light reflected from the optic was then incident normal to the RC flat mirror and reflected back through the fiber. This ensured that the incident beam was aligned to the angle of the RC eigenmode (assuming no changes in the alignment of the mirror after the science campaign) to better than 1 mrad. The results are shown in Table 5. Similar results were expected for the pair of light tight housing mirrors and the pair of Mach-Zehnder mirrors, and the measurements of LT1 and LT2 both resulted in a transmissivity of $12.4\pm 0.5\rm\,ppm$. MZ3 was measured to have a transmissivity of $452\pm 10\rm\,ppm$ while the transmissivity of MZ4 was $275\pm 10\rm\,ppm$. While MZ3 and MZ4 were expected to have the same transmissivity as their dielectric coatings were applied during the same coating run, the discrepancy between them is believed to be related to differences in their angles of incidence with respect to the optical axis of the system. The uncertainty in these measurements is due to a combination of the uncertainty in the calibration of the power meters used in the setup, the presence of stray light at the power meters measuring the weak transmitted signals, and the uncertainty of the angle of incidence of the laser on the mirrors.
In-situ measurements of the on-axis COB optics were also performed in between $\rm S_{{}_{\perp}}$ and $\rm S_{{}_{\parallel}}$. Here LT1 was measured using a power meter to measure the HPL directly at the input window to the vacuum system in NL and then at the High-Power port of the CH vacuum tank (see Figure 6). This gave a value of $11\pm 1$ ppm. LT2 was measured with a similar method, but instead measuring the RL power just before the Injection Port at the CH tank, also giving a value of $11\pm 1$ ppm. The combined transmissivity of MZ3 and MZ4 was measured using a power meter to measure the HPL power at the PLL2 port, and then at the Mach-Zehnder port by measuring the amplitude of the RL-HPL beatnote. The spatial overlap of the RL-HPL beatnote at optimal alignment could be directly measured at the PLL2 port because it was possible to measure the static power of both lasers there. This value could then be used to compensate for these effects in the measurement of the beatnote amplitude at the Mach-Zehnder port. [^29] It should be noted that to find the beatnote amplitude at each port for the ‘optimal’ alignment, RL was realigned to maximize the the beatnote amplitude in both cases. This was necessary to compensate for the alignment drift of the MZ on the COB. The combined transmissivity of the MZ mirrors found in this measurement was $0.137\pm 0.007$ ppm.
The result of the combined transmissivities of the MZ mirror from the ex-situ measurements gave $0.124\pm 0.005$ ppm. Thus, it is clear that the individual measurements of each of the on-axis COB optics agree between the two methods. Even more relevant is the total transmissivity $T_{{}_{\rm COB}}$, in which a value of $(1.9\pm 0.1)\times 10^{-17}$ was found for the ex-situ measurement with $(1.7\pm 0.2)\times 10^{-17}$ measured for the in-situ measurements. In this case the error bars of the measurements do show some overlap and the ex-situ measurements nearly lie within the error bars of the in-situ measurements.
Finally, the most relevant parameter for the calibration of the results in terms of $\mathcal{P}_{\gamma\leftrightarrow\phi}$ is the product of the transmissivity of COB optics and the cavity mirror RC2. For this, the ex-situ measurements gave $T_{{}_{\rm COB}}T_{{}_{\rm RC2}}=(9.7\pm 1.2)\times 10^{-23}$ while the in-situ results were $T_{{}_{\rm COB}}T_{{}_{\rm RC2}}=(8.5\pm 1.4)\times 10^{-23}$. Here in both cases, the measured values lie within the other measurements error bars. This gives us confidence in the accuracy of the overall calibration of the results. We consider the method for the ex-situ measurements to be more robust as we were able to directly access the optics. Therefore, the ex-situ measurements were used for the actual calibration of the results.
IV.2.3 Polarization Overlap
The error in the orientation of the optical tables and magnetic fields is on the order of 100 µrad. The polarization state of the HPL for $\rm S_{{}_{\perp}}$ was tuned for maximum transmission through LT1 (which was rigidly mounted to the COB). For $\rm S_{{}_{\parallel}}$ an additional $\lambda/2$ wave-plate, whose orientation was precisely aligned, was also located on the COB in the path to LT1. The polarization state of the AL and LO are set by a rotatable half wave-plate in NR. The exact angle of the ‘roll’ of the COB with respect to the optical axis was not measured.
IV.2.4 Field Overlap
In the absence of significant effects that can alter the amplitude or alignment of the HPL as it travels between the production and regeneration areas, the coupling measured during the open shutter periods should be representative of the spatial coupling of the BSM field to the RC. The spatial component of the field overlap between the HPL and the RC measured by the science detector was expressed earlier as Equation 5 in Section I.1. The square of the spatial field overlap, $\eta_{{xy},j}^{2}$, represents the spatial overlap in terms of power, and in Equation 76, this is expressed as a function of measured parameters. It can be found by calculating the ratio of the measured power during an open shutter period $P_{{\rm{open}},j}$ versus the expected power, given by the product of the power of the laser traveling through the production magnet string $P_{\rm i}$, the transmissivity of the on-axis COB optics $T_{{}_{\rm COB}}$, and the transmissivity of the cavity itself for the HPL frequency, $T_{{}_{\rm RC}}$, as shown by
$\eta_{{xy},j}^{2}$ $=\frac{P_{{\rm open},j}}{T_{\!{}_{\rm COB}}T_{\!{}_{\rm RC}}P_{\rm i}}$ (75)
$=\frac{|C_{{\rm open},j}|^{2}}{T_{\!{}_{\rm COB}}T_{\!{}_{\rm RC}}P_{\!{}_{\rm LO}}G_{\!{}_{\rm AC}}^{2}T_{\!{}_{\rm sci}}\eta_{\!{}_{xy,{\rm LO}}}^{2}}.$ (76)
Here, by using $T_{\rm RC}$ and not $T_{\rm RC,max}$, $\eta_{{xy},j}^{2}$ is defined as the spatial overlap rather than the total field overlap, as the component of the longitudinal overlap in $P_{{\rm open},j}$ is removed by dividing by $T_{\rm RC}$.
(a) $\rm S_{{}_{\perp}}$
(b) $\rm S_{{}_{\parallel}}$
(c) $\rm S_{{}_{\perp}}$
(d) $\rm S_{{}_{\parallel}}$
Figure 21: The square of the spatial component of the field overlap $\eta_{xy}^{2}$ and $\eta_{xy,{\rm veto}}^{2}$ measured at the science detector and veto detectors respectively, from data taken during the maintenance periods of $\rm S_{{}_{\perp}}$ is shown in (a) and from $\rm S_{{}_{\parallel}}$ in (b). Here the black lines in (a) and (b) give any times at which the cavity alignment was actively changed, either to optimize the spatial field overlap or the resonant enhancement. The square of the total field overlap measured at the science detector $\eta^{2}$ and the veto detector $\eta_{\rm veto}^{2}$ is shown for $\rm S_{{}_{\perp}}$ in (c) and $S_{{}_{\parallel}}$ in (d). The black lines in (c) and (d) again shown any times at which the cavity was realigned, as well as times in which the length of the cavity was change to optimize the spectral component of the field overlap. In the case of (c) the length of the cavity was never changed during $\rm S_{{}_{\perp}}$.
The results of measuring $\eta_{{xy},j}^{2}$ are shown in Figures 20(a) and 20(b) respectively for $\rm S_{{}_{\perp}}$ and $\rm S_{{}_{\parallel}}$. For $\rm S_{{}_{\perp}}$, the average spatial overlap between the HPL and the RC during the closed shutter periods was estimated to be 47% with a range of 41% to 54% and RMS fluctuations from the average of 4%. The uncertainty in these measurements varied from 7% to 10%, with an average value of 9%, and was due mostly to the uncertainties in the measured transmissivities of the on-axis COB mirrors and the flat cavity mirror mounted to the COB (RC2), with some additional contribution due to the uncertainty on the calibration series. During the closed shutter periods of $\rm S_{{}_{\parallel}}$ the mean value of $\eta_{{xy}}^{2}$ was estimated to be 48% with a minimum of 42%, a maximum of 60%, and an RMS of 6%. On these measurements, the uncertainty was again driven by the uncertainties on the mirror transmissivities like $\rm S_{{}_{\perp}}$, along some additional uncertainty due to the calibration series. For $\rm S_{{}_{\parallel}}$ the uncertainty on $\eta_{xy}^{2}$ varied between 6% and 11% with an average value of 9%.
$\eta^{2}$ $=\eta_{{xy},j}^{2}\eta_{\hat{z}}^{2}$ (77)
$=\frac{|C_{{\rm open},j}|^{2}}{T_{{}_{\rm COB}}T_{\rm RC,max}P_{\!{}_{\rm LO}}G_{\!{}_{\rm AC}}^{2}T_{\!{}_{\rm sci}}\eta_{\!{}_{xy,{\rm LO}}}^{2}}.$ (78)
With $\eta^{2}$, the expected power exiting the RC due to the reconversion of the BSM field to an electromagnetic field can be calculated.
In this case $T_{\rm RC,max}$ is used, so the effects of the longitudinal overlap are included in this calculation.
Measurements at the science detector of the total overlap $\eta^{2}$ are shown for $\rm S_{{}_{\perp}}$ and $\rm S_{{}_{\parallel}}$ in Figures 20(c) and 20(d). During the closed shutter periods of $\rm S_{{}_{\perp}}$, the mean value of the total overlap was estimated to be 43% with a range from 33% to 51%, RMS fluctuations from the mean of 4%, with an uncertainty that varied between 6% and 10% and had an average value of 8%. These measurements were limited by the same sources of uncertainty as $\eta_{{xy},j}^{2}$. For $\rm S_{{}_{\parallel}}$, the average estimated value of $\eta^{2}$ at the science detector during the closed shutter periods was 41% with a range from 21% to 58%, an RMS of 9%, and an uncertainty between 4% and 10% with an average value of 8%. The uncertainty on these measurements was also due to the same effects as were just mentioned for $\rm S_{{}_{\perp}}$.
IV.2.5 Optimizing the Field Overlap
IV.3 Veto Detection System
The veto detector can also be used to measure the overlap between the HPL and the RC. One advantage of using the veto detector is that the HPL does not need to pass through the cavity to reach it and therefore the HPL power is higher at the veto PD. One complication arises from the fact that the optical losses on the path from the output mirror of the RC to the veto PD can not be directly measured, as there are a significant number of optics in the portion of this path inside the central vacuum chamber. These losses can be calculated using the ratio of the power of the AL measured at the veto PD to the expected power of AL given its input power and coupling efficiency to the RC, as well as the transmissivity of the RC. The uncertainties related to the transmissivity of the cavity therefore still play a role in the power overlap measured at the veto detector. Section IV.3.1 explains how the HPL power at the veto detector is calibrated, describes the measurements of the gain of the veto detection system, and presents measurements of the AL power at the veto PD. Section IV.3.2 goes into detail on the calibration of the optical losses between the cavity and the veto detector. Finally, Section IV.3.3 explains how the calculation of the power overlap was performed and summarizes the results of these measurements during first science campaign.
IV.3.1 Signal Power
@@
The open shutter series is defined for the veto detection system in the same way as for the science detector and will be referred to as $C^{\rm veto}_{{\rm open},j}$. Like for the science detection system, $C^{\rm veto}_{{\rm open},j}$ is proportional to the coherent signal power to HPL present at the veto detector and is in units of $\rm V^{2}/W$.
It is useful to define the optical power present at the veto PD, $P^{\rm veto}_{{\rm open},j}$, in terms of $C^{\rm veto}_{{\rm open},j}$:
$P^{\rm veto}_{{\rm open},j}=\frac{P_{\rm i}|C^{\rm veto}_{{\rm open},j}|^{2}}{P_{\!{}_{\rm AL,veto}}G_{\!{}_{\rm AC,veto}}^{2}T_{\!{}_{\rm veto}}}.$ (79)
@@
The power measured at the veto photodetector $P_{\rm AL,veto}$ is given by the DC voltage output by the veto PD $V_{\rm DC,veto}$, divided by its DC conversion gain $G_{\rm DC,veto}$:
@@
This is an important parameter as it plays a direct role in the calibration of $C^{\rm veto}_{{\rm open},j}$ in terms of optical power. It can also be used to estimate the optical losses in the path from the RC output mirror to the veto PD discussed in the following section. The DC conversion gain $G_{\!{}_{\rm DC,veto}}$ of the veto detection system was measured by varying the power incident on the veto detector, and measuring the voltage at the monitoring point used during the maintenance periods. This gave a value of $G_{\!{}_{\rm DC,veto}}=86\pm 1$ V/mW. The error in this measurement is due to the variance in the results over multiple measurements. The primary reason for the difference between the AC and DC gain, was the presence of a three-way splitter in the AC path that allowed the veto PD to also be used as the sensor for PLL1.
IV.3.2 Optical Losses on Veto Path
The transmissivity of the optics between the cavity and the veto PD, denoted by $T_{\rm veto}$, could not be directly measured because the RC transmitted power does not directly leave the CH vacuum tank like the cavity reflected power in NR. Instead, it had to be estimated during the runs using parameters measured during the maintenance periods. One of these parameters is the fraction of power of the AL that directly couples to the RC, referred to in the following equation as $\Delta P_{\rm AL,NR}$. [^30] It should be pointed out that this parameter is different from $P_{\rm AL,veto}$ described in the previous section. This is measured by dividing the drop in the DC voltage output at the science PD $\Delta V_{\rm AL,sci}$, when AL is frequency stabilized to the cavity, by the DC gain of the science PD detection chain ($G_{\rm DC,sci}$) and by $T_{\rm sci}$ as expressed by
@@ IV.3.2 Optical Losses on Veto Path
IV.3.3 Field Overlap
@@
To calculate the HPL-RC spatial power overlap (the absolute square of $\eta_{xy}$ given in Equation 5) measured by the veto detection system $\eta_{xy,{\rm veto}}^{2}$, the open shutter power defined in Section IV.3.1 is divided by the expected power of the HPL passing through the production magnet string and the COB optics, along with the reflectivity of the cavity from the direction of CH:
$\eta_{xy,{\rm veto}}^{2}=\frac{P^{\rm veto}_{{\rm open},j}}{R_{\rm RC,CH}T_{{}_{\rm COB}}P_{\rm i}}.$ (86)
@@
When expressed in terms of measured parameters, the spatial overlap measured by the veto detection system $\eta_{xy,{\rm veto}}$ and the corresponding total overlap $\eta_{\rm veto}$ (the product of the spatial overlap with the longitudinal overlap) are
$\eta_{xy,{\rm veto}}^{2}\simeq\frac{|C^{\rm veto}_{{\rm open},j}|^{2}\Delta V_{\rm AL}G^{2}_{\rm DC,veto}T_{\!{}_{\rm RC2}}}{G_{\!{}_{\rm AC,veto}}^{2}T_{\rm sci}\left|1-\frac{T_{\!{}_{\rm RC2}}f_{0}\tau}{1-i2\pi\tau\Delta\nu}\right|^{2}T_{{}_{\rm COB}}V^{2}_{\rm DC,veto}G_{\rm DC,sci}\left(\frac{2}{f_{0}\tau}-T_{\!{}_{\rm RC1}}\right)}{,\hskip 14.22636pt\rm and}$ (87)
$\eta_{\rm veto}^{2}\simeq\frac{|C^{\rm veto}_{{\rm open},j}|^{2}\Delta V_{\rm AL}G^{2}_{\rm DC,veto}T_{\!{}_{\rm RC2}}}{G_{\!{}_{\rm AC,veto}}^{2}T_{\rm sci}\left((1-T_{\!{}_{\rm RC2}}f_{0}\tau)^{2}+(2\pi\tau\Delta\nu)^{2}\right)T_{{}_{\rm COB}}V^{2}_{\rm DC,veto}G_{\rm DC,sci}\left(\frac{2}{f_{0}\tau}-T_{\!{}_{\rm RC1}}\right)}.$ (88)
@@
Equations 69, 79, 80, 85, and 63 are used to derive these expressions. Figures 20(a) and 20(b) show the HPL-RC spatial overlap measured at the veto detector during the maintenance periods of the science runs as the orange data points. During the closed shutter periods of $\rm S_{{}_{\perp}}$, the estimated spatial overlap had an average value of 53% with a range of 40% to 63%, RMS fluctuations from the mean of 6%, and an uncertainty of between 6% and 10% with an average uncertainty of 8%. For $\rm S_{{}_{\parallel}}$ the average value was 49% with a minimum of 34%, a maximum value of 66%, an RMS of 8%, and an uncertainty between 5% and 10% with an average value of 7%. For both runs the uncertainty was primarily limited by the uncertainty on the measured transmissivities of the COB optics and mirror RC2. While RC2 does not play a direct role in the HPL power measured at the veto PD, its uncertainty does effect the measurement of $T_{\rm veto}$.

### v2

@@ IV.2.3 Open Shutter Power Ratio
To find $T_{\!{}_{\rm sci}}$, the ratio of power incident on the RC cavity mirror versus the power incident on the science detector was calculated. This was done several times before, during, and after the first science campaign and gave values between 80% and 90%. The value used to calculate the field overlap was $85\%\pm 5\%$. The losses were mostly due to optical pickoffs used to sample the light reflected from the cavity for the PDH frequency stabilization system as well as other diagnostic purposes. Changes in the total reflectivity of the input optics that appear to be humidity dependent were also observed, along with clipping losses on the free aperture of the input optics after the beam expanding telescope in NR. Both of these effects played a role in the uncertainty of the results of $T_{\!{}_{\rm sci}}$.
@@
The open shutter calibration series for the veto detection system $C^{\rm veto}_{{\rm open}}$, is defined in the same way as for the science detector and is also in units of $\rm V^{2}/W$. The power ratio $P^{\rm veto}_{{\rm open}}/P_{\rm i}$ for the veto detector can be expressed as
$\frac{P^{\rm veto}_{{\rm open}}}{P_{\rm i}}=\frac{|C^{\rm veto}_{{\rm open}}|^{2}}{P_{\!{}_{\rm AL,veto}}G_{\!{}_{\rm AC,veto}}^{2}T_{\!{}_{\rm veto}}}.$ (70)
@@
The power measured at the veto photodetector $P_{\rm AL,veto}$ is an important parameter as it plays a direct role in the calibration of $C^{\rm veto}_{{\rm open}}$ in terms of optical power. It can also be used to estimate the optical losses in the path from the RC output mirror to the veto PD discussed in the following section. $P_{\rm AL,veto}$ can be calculated by dividing the DC voltage output by the veto PD $V_{\rm DC,veto}$, by its DC conversion gain $G_{\rm DC,veto}$:
@@
The DC conversion gain $G_{\!{}_{\rm DC,veto}}$ of the veto detection system was measured by varying the power incident on the veto detector, and measuring the voltage at the monitoring point used during the maintenance periods. This gave a value of $G_{\!{}_{\rm DC,veto}}=86\pm 1$ V/mW. The error in this measurement is due to the variance in the results over multiple measurements. The primary reason for the difference between the AC and DC gain, was the presence of a three-way splitter in the AC path that allowed the veto PD to also be used as the sensor for PLL1.
One complication in the measurement of $P^{\rm veto}_{{\rm open}}/P_{\rm i}$ arises from the fact that the transmissivity of the optics on the path from the output mirror of the RC to the veto PD, $T_{\rm veto}$, can not be directly measured, as there are a significant number of optics in the portion of this path inside the central vacuum chamber. $T_{\rm veto}$ can be calculated using the ratio of the power of the AL measured at the veto PD to the expected power of AL given its input power and coupling efficiency to the RC, as well as the transmissivity of the RC. This is measured by dividing the drop in the DC voltage output at the science PD $\Delta V_{\rm AL,sci}$, when AL is frequency stabilized to the cavity, by the DC gain of the science PD detection chain ($G_{\rm DC,sci}$) and by $T_{\rm sci}$ as expressed by
@@
IV.2.4 Spatial Overlap Results
The spatial coupling of the HPL to the RC measured during the open shutter periods should be representative of the coupling of the BSM field as long as there are no significant effects that can alter the amplitude or alignment of the HPL as it travels between the production and regeneration areas. The spatial component of the field overlap between the HPL and the RC measured by the science detector was expressed earlier as Equation 5 in Section I.1. The square of the spatial field overlap, $\eta_{{xy}}^{2}$, represents the spatial overlap in terms of power and can be expressed as a function of measured parameters as:
$\eta_{{xy}}^{2}=\frac{|C_{{\rm open}}|^{2}}{T_{\!{}_{\rm COB}}T_{\!{}_{\rm RC}}P_{\!{}_{\rm LO}}G_{\!{}_{\rm AC}}^{2}T_{\!{}_{\rm sci}}\eta_{\!{}_{xy,{\rm LO}}}^{2}}.$ (77)
Here, $T_{\rm RC}$ is used instead of $T_{\rm RC,max}$ because $\eta_{{xy}}^{2}$ is defined as the spatial overlap rather than the total field overlap.
(a) $\rm S_{{}_{\perp}}$
(b) $\rm S_{{}_{\parallel}}$
(c) $\rm S_{{}_{\perp}}$
(d) $\rm S_{{}_{\parallel}}$
Figure 21: The field overlap between the HPL and RC measured at the science detector and veto detectors respectively, from data taken during the maintenance periods of $\rm S_{{}_{\perp}}$ and $\rm S_{{}_{\parallel}}$. The squared spatial overlap $\eta_{xy}^{2}$ is show in (a) for $\rm S_{{}_{\perp}}$ and (b) for $\rm S_{{}_{\parallel}}$. The squared total field overlap $\eta^{2}$ is show in (c) for $\rm S_{{}_{\perp}}$ and (d) for $\rm S_{{}_{\parallel}}$. In all four plots the black lines indicate any times at which the cavity alignment was actively changed, either to optimize the spatial field overlap or the resonant enhancement, while in (d) they also indicate times in which the length of the cavity was change to optimize the spectral component of the field overlap. In the case of (c) the length of the cavity was never changed during $\rm S_{{}_{\perp}}$.
The results of measuring $\eta_{{xy}}^{2}$ are shown in Figures 20(a) and 20(b) respectively for $\rm S_{{}_{\perp}}$ and $\rm S_{{}_{\parallel}}$. For $\rm S_{{}_{\perp}}$, the average spatial overlap between the HPL and the RC during the closed shutter periods was estimated to be 47% with a range of 41% to 54% and RMS fluctuations from the average of 4%. The uncertainty in these measurements varied from 7% to 10%, with an average value of 9%, and was due mostly to the uncertainties in the measured transmissivities of the on-axis COB mirrors and the flat cavity mirror mounted to the COB (RC2), with some additional contribution due to the uncertainty on the calibration series. During the closed shutter periods of $\rm S_{{}_{\parallel}}$ the mean value of $\eta_{{xy}}^{2}$ was estimated to be 48% with a minimum of 42%, a maximum of 60%, and an RMS of 6%. The uncertainty of these measurements was again driven by the uncertainties on the mirror transmissivities like in $\rm S_{{}_{\perp}}$, along with some additional uncertainty due to the calibration series. For $\rm S_{{}_{\parallel}}$ the uncertainty on $\eta_{xy}^{2}$ varied between 6% and 11% with an average value of 9%.
@@ IV.2.4 Spatial Overlap Results
To calculate the HPL-RC spatial power overlap measured by the veto detection system $\eta_{xy,{\rm veto}}^{2}$, the open shutter power defined in Section IV.2.3 is divided by the expected power of the HPL passing through the production magnet string and the COB optics, along with the reflectivity of the cavity from the direction of CH:
$\eta_{xy,{\rm veto}}^{2}=\frac{P^{\rm veto}_{{\rm open}}}{R_{\rm RC,CH}T_{{}_{\rm COB}}P_{\rm i}}.$ (78)
@@
The spatial overlap measured by the veto detection system can be expressed in terms of measured parameters as:
$\eta_{xy,{\rm veto}}^{2}\simeq\frac{|C^{\rm veto}_{{\rm open}}|^{2}\Delta V_{\rm AL}G^{2}_{\rm DC,veto}T_{\!{}_{\rm RC2}}\left(1+(2\pi\tau\Delta\nu)^{2}\right)}{G_{\!{}_{\rm AC,veto}}^{2}T_{\rm sci}\left((1-T_{\!{}_{\rm RC2}}f_{0}\tau)^{2}+(2\pi\tau\Delta\nu)^{2}\right)T_{{}_{\rm COB}}V^{2}_{\rm DC,veto}G_{\rm DC,sci}\left(\frac{2}{f_{0}\tau}-T_{\!{}_{\rm RC1}}\right)}.$ (79)
@@
Equations 58, 62, 70, 71, and 76 are used to derive these expressions. Figures 20(a) and 20(b) show the HPL-RC spatial overlap measured at the veto detector during the maintenance periods of the science runs as the orange data points and the results of $\eta_{xy,{\rm veto}}^{2}$ agree very well witht he data acquire at the science detector. During the closed shutter periods of $\rm S_{{}_{\perp}}$, the estimated spatial overlap had an average value of 53% with a range of 40% to 63%, RMS fluctuations from the mean of 6%, and an uncertainty of between 6% and 10% with an average uncertainty of 8%. For $\rm S_{{}_{\parallel}}$ the average value was 49% with a minimum of 34%, a maximum value of 66%, an RMS of 8%, and an uncertainty between 5% and 10% with an average value of 7%. For both runs the uncertainty was primarily limited by the uncertainty on the measured transmissivities of the COB optics and mirror RC2. While RC2 does not play a direct role in the HPL power measured at the veto PD, its uncertainty does effect the measurement of $T_{\rm veto}$.
IV.2.5 Field Overlap Results
$\eta^{2}$ $=\eta_{{xy}}^{2}\eta_{\hat{z}}^{2}$ (80)
$=\frac{|C_{{\rm open}}|^{2}}{T_{{}_{\rm COB}}T_{\rm RC,max}P_{\!{}_{\rm LO}}G_{\!{}_{\rm AC}}^{2}T_{\!{}_{\rm sci}}\eta_{\!{}_{xy,{\rm LO}}}^{2}}.$ (81)
In this case $T_{\rm RC,max}$ is used because the effects of the longitudinal overlap must be included in this calculation.
Measurements at the science detector of the total overlap $\eta^{2}$ are shown for $\rm S_{{}_{\perp}}$ and $\rm S_{{}_{\parallel}}$ in Figures 20(c) and 20(d). During the closed shutter periods of $\rm S_{{}_{\perp}}$, the mean value of the total overlap was estimated to be 43% with a range from 33% to 51%, RMS fluctuations from the mean of 4%, with an uncertainty that varied between 6% and 10% and had an average value of 8%. These measurements were limited by the same sources of uncertainty as $\eta_{{xy}}^{2}$. For $\rm S_{{}_{\parallel}}$, the average estimated value of $\eta^{2}$ at the science detector during the closed shutter periods was 41% with a range from 21% to 58%, an RMS of 9%, and an uncertainty between 4% and 10% with an average value of 8%. The uncertainty on these measurements was also due to the same effects as were just mentioned for $\rm S_{{}_{\perp}}$.
Using Equation 79 to express $\eta_{xy,{\rm veto}}^{2}$, the total field overlap between the HPL and the RC can be expressed in terms of measured parameters as
$\eta_{\rm veto}^{2}\simeq\frac{\eta_{xy,{\rm veto}}^{2}}{\eta_{\hat{z}}^{2}}\simeq\frac{\eta_{xy,{\rm veto}}^{2}}{1+(2\pi\tau\Delta\nu)^{2}}.$ (82)